Calculate the pH of a solution that is 0.20M with benzoic acid and .18with its salt, sodium benzoate


1
Expert's answer
2022-01-17T17:04:03-0500


pH"=" pk"_a" "+log\frac{[A^-]}{[HA]}"


k"_a=6.5\u00d710^{-5}"


pH"=-log(6.5×10^{-5})+log\frac{0.18}{0.20}"


"=4.1871-0.04576"

"=4.1413"